OpenStudy (anonymous):

the width of a rectangle 8in. greater than the length. the perimeter is 102in. find the dimensions of the rectangle. linear equations

OpenStudy (anonymous):

@mathstudent55 @misty1212

OpenStudy (misty1212):

HI!! 

OpenStudy (anonymous):

HI. :)

OpenStudy (misty1212):

lets add up to find the perimeter

OpenStudy (misty1212):

\[x+x+8+x+x+8=4x+16\] right?

OpenStudy (anonymous):

right

OpenStudy (misty1212):

and you know that is (102\) so your job is to solve \[4x+16=102\] for \(x\)

OpenStudy (misty1212):

can you do that?

OpenStudy (anonymous):

yes

OpenStudy (misty1212):

ok let me know what you get

OpenStudy (anonymous):

kk one sec

OpenStudy (anonymous):

-26

OpenStudy (anonymous):

@misty1212

OpenStudy (misty1212):

oh no

OpenStudy (misty1212):

\[4x+16=102\] first subtract \(16\) from both sides

OpenStudy (anonymous):

-86

OpenStudy (anonymous):

@misty1212

OpenStudy (mathstudent55):

You did 16 - 102. You need to do 102 - 16

OpenStudy (anonymous):

ohhhhh

OpenStudy (anonymous):

86-

OpenStudy (anonymous):

@misty1212 @mathstudent55

OpenStudy (mathstudent55):

Good, Now you have 4x = 86

OpenStudy (anonymous):

21.5

OpenStudy (misty1212):

lol what does \(86-\) mean ??

OpenStudy (mathstudent55):

Correct.

OpenStudy (anonymous):

thanks. :)

OpenStudy (mathstudent55):

That is x. Now you need x + 8 = 21.5 + 8 What is 21.5 + 8 ?

OpenStudy (anonymous):

one sec

OpenStudy (anonymous):

29.5

OpenStudy (mathstudent55):

Correct.

OpenStudy (anonymous):

thanks.